# Excitation of Inertial Modes in 3D Simulations of Rotating Convection

### v1

@@ 3.2 Frequency Spectra
Figure 4 shows the frequency spectra of the kinetic energy at $r=0.85r_{o}$, for all the simulations with the exception of the non-rotating case, which is omitted as it shows no notable spectral features. Distinct coherent peaks in the spectrum reveal the excitation of modes, and the broader structure reflects the background of convective turbulence. The total frequency spectra refer to the ones summed over all $\ell$ and $m$, while ones for specific $m$ are summed over all $\ell$ for that specific order.
Note that each spectrum contains an $m=0$ contribution that is much larger at very low frequencies. This corresponds to the “zero-frequency” mean flows associated with differential rotation, which evolve on a timescale comparable to the Ekman spin-up time, $\tau_{\mathrm{EK}}\sim(1/\mathrm{Ek})^{1/2}\Omega_{0}^{-1}$, with the corresponding frequency range marked by the vertical dashed lines in each panel.
@@
As the convective Rossby number decreases to 0.35 (panel d), a prominent $m=2$ mode appears, accompanied by a weaker $m=4$ mode whose frequency is twice that of $m=2$, indicating a non-linear self-interaction that resemble the ones described in Barik2018 for the spherical Couette system. At $\mathrm{Ro_{c}}\approx 0.26$ (panel e), the spectrum is dominated by an $m=3$ mode, followed by an $m=2$ signal at a nearby frequency, and low amplitude $m=1$ and $m=6$ modes. The frequency of the $m=6$ mode is exactly twice that of the $m=3$ mode, indicating that it arises from a self-interaction of the lower mode. At the lowest Rossby number in the suite, $\mathrm{Ro_{c}}\approx 0.17$ (panel f), we find three small-amplitude modes corresponding to $m=1$, $m=2$, and $m=6$, with the $m=2$ mode having a frequency roughly twice that of the $m=1$ mode. We emphasize that the amplitude of the excited modes depends on both the damping and the excitation mechanisms, which in our simulations are likely dominated by viscosity and the degree of differential rotation. Although rotation weakens convection, particularly at low Rossby numbers, all of our rotating simulations exhibit strong differential rotation. However, we do not find any clear correlation between the mode amplitude and the strength of the differential rotation.
For comparison, each panel also shows the frequency of the spectrum normalized to the convective frequency $\omega_{\rm conv}\sim R/U_{\rm conv}$, where $U_{\rm conv}$ is the standard non-rotating convective speed from mixing-length theory. We can see that the power spectrum is fairly flat for $\omega/\omega_{\rm conv}\lesssim 1$, and it generally falls steeply for $\omega/\omega_{\rm conv}\gtrsim 1$, as expected. However, it does not clearly exhibit the $\mathcal{P}(\omega)\propto\omega^{-2}$ spectrum expected for Kolmogorov turbulence with $\omega>\omega_{\rm conv}$ (goldreich:90; goldreich:94). We might expect different slopes in three regimes, with the lowest frequencies being $\omega<\omega_{\rm conv}$, the intermediate frequency regime in $\omega_{\rm conv}<\omega<2\Omega_{0}$, and the high-frequency regime being $\omega>2\Omega_{0}$. Our $\mathrm{Ro_{c}}\approx 0.17$ simulation appears to have a shallower slope to the power spectrum in the intermediate frequency regime than the high-frequency regime. For the other simulations, the separation in scales between $\omega_{\rm conv}$ and $2\Omega$ is too small to see any clear differences.
@@ 3.3 Mode Eigenfunctions
Figure 6 shows the eigenfunctions of several modes excited in the simulations. We find that $u_{\phi}$ provides the strongest contribution to the power spectrum in these modes. Thus, we filter the zonal component of the velocity in space and time, isolating the azimuthal wavenumber $m$ and the mode frequency. The presence of differential rotation and contamination from convective motions makes the eigenfunction shapes complex and difficult to interpret. Nevertheless, these projections confirm that all the inertial modes in our simulations, except for the $m=0$ mode at $\mathrm{Ro_{c}}\approx 0.53$, exhibit symmetry about the equator, as also suggested by the dominance of the symmetric component in Figure 5.
3.4 Internal Shear Layers and Mode Attractors
Another useful measure is the meridional distribution of the zonal kinetic energy, which reveals the presence of internal shear layers (StewartsonRickard1969; Kerswell1995; Rieutord2001). These features originate at the critical co-latitudes defined by $\cos\theta_{c}=\omega/2\Omega_{0}$, forming rays tilted by an angle $\pi/2-\theta_{c}$ with respect to the rotation axis. They trace the characteristics of the hyperbolic Poincaré equation that governs inertial modes. Under realistic boundary conditions (no-slip or stress-free), the corresponding boundary-value problem becomes ill-posed, producing singularities along these paths (BondiLyttleton1953; StewartsonRickard1969; Kerswell1995). Viscosity removes these singularities and produces narrow regions of strong shear along the attractor paths.
@@ 3.5 Low Prandtl number
We find several differences relative to the model using $\mathrm{Pr}=1$. For example, because the effective buoyancy scales as $\mathrm{Ra}/\mathrm{Pr}$, the flow velocities are much larger in the $\mathrm{Pr}=0.1$ simulation (compare the last two rows of Table 1). Another difference is the differential rotation. While the $\mathrm{Pr}=1$ case shows weak flows at high latitudes (see panels a of Figures 1 and 2), the $\mathrm{Pr}=0.1$ model exhibits vigorous flows at all latitudes (see Figure 8).
The frequency spectrum of the modes also shows remarkable differences with respect to the $\mathrm{Pr}=1$ case. As shown by Figure 9, mode excitation is more efficient at low Pr, with several large peaks appearing over the range m=1m=1–8. These inertial modes are mostly equatorially symmetric and drift retrograde in the rotating frame. In contrast, the $\mathrm{Pr}=1$ case exhibits much weaker excitation, with only three modes of very small amplitude (see panel f of Figure 4), even though the convective Rossby number $\mathrm{Ro_{c}}\approx 0.17$ remains the same in both simulations. As discussed previously, the dominant $m=1$ mode in this case is equatorially symmetric but prograde in the rotating frame. Finally, another difference is the spatial location of the modes. While the low amplitude $m=1$ mode observed for $\mathrm{Pr}=1$ is confined to the equatorial region (see Figure 6), the modes for $\mathrm{Pr}=0.1$ (including the $m=1$ mode) have most of their power from mid to high latitudes (see Figure 10).
4 Summary and Discussion
@@
In the spatial domain, the kinetic energy power spectra of the flow resemble those reported in previous numerical simulations of rotating convection (Featherstone2016; Hindman2020), with energy dominated by axisymmetric motions that produce strong peaks at odd spherical harmonic degrees $\ell$ (see Figure 3). Interestingly, the zonal component of the kinetic energy power spectra exhibits an $\ell^{-5}$ scaling, as observed in the spectra of the zonal flows of the gas and ice giants (e.g., Sanchez2000; Sukoriansky2002), and predicted analytically for zonostrophic turbulence (see the recent preprint by Boning2024; Cabanes2024).
In the frequency domain, the low frequency range of the power spectra, which is limited by the frequency associated with the Ekman spin-up time of the flow $\tau_{\rm Ek}\sim 1/\sqrt{\rm Ek}\Omega$, is dominated by the $m=0$ contribution of the differential rotation. At higher frequencies, the spectra show an overall flat distribution with superimposed peaks, followed by a rapid decay for $\omega/\omega_{\rm conv}>1$, i.e., for frequencies larger than the convective frequency (see Figures 4 and 9). These coherent peaks, which correspond to inertial modes excited within the flow, are only observed in cases with $\mathrm{Ro_{c}}\lesssim 0.53$, while they are absent at larger Rossby numbers. This is expected since the amount of differential rotation increases for more rapidly rotating flows. However, the amplitude of the peaks does not appear to correlate with the Rossby number and depends on the growth rate and saturation of each mode, influenced by both the damping and excitation mechanisms. The Prandtl number nevertheless appears to play a more important role in mode excitation, since the $\mathrm{Pr}=0.1$ simulation exhibits a richer and more vigorous spectrum of inertial modes than any of the simulations with $\mathrm{Pr}=1$.
For the most part, the inertial modes in our simulations are nonaxisymmetric, retrograde in the rotating frame, symmetric about the equator, and span azimuthal orders $1\leq m\leq 8$, with frequencies $\omega/2\Omega_{0}\lesssim 0.2$ (see Figure 5 and 9). Their eigenfunctions are complex, likely due to the effects of turbulent convection and differential rotation, and show that nearly all modes are confined to mid and high latitudes. This suggests that their excitation may be linked to shear instabilities within the differentially rotating convection zone. Further, their zonal energy distributions reveal internal shear layers that resemble mode attractors, particularly at high latitudes (see Figure 7).
@@
There are several differences between our results and those in previous work. Zhang2001 found a special class of equatorially symmetric inertial modes which strongly satisfy the Taylor-Proudman constraint (i.e., if the pressure gradient and the Coriolis force balance each other, see Proudman1916; Taylor1917). In such scenario, the modes are symmetric with respect to the equator, their frequencies satisfy $i\omega~{\mbox{{$u$}}}=-\nabla P/\rho-2\Omega_{0}{\mbox{{$\hat{z}$}}}\times{\mbox{{$u$}}}\approx 0$, and in turn, they are expected to be nearly geostrophic ($\omega/\Omega_{0}\ll 1$). Those modes are referred to as “slow” inertial modes (Barik2018). In contrast, equatorially antisymmetric modes do not satisfy the Taylor-Proudman constraint, and can have much larger frequencies than the slow inertial modes (see, also Wicht2014). The modes we observe in our simulations are mainly equatorially symmetric and all have frequencies $\omega/2\Omega_{0}\sim 0.1$. However, the equatorially symmetric modes are not close to being geostrophic.
Another difference is about the excitation mechanism. In general, the inertial modes observed in simulations and laboratory experiments of spherical Couette flows have been attributed to shear instabilities. It is likely that many of the modes in our simulations are excited simultaneously by both convection and differential rotation, with only a few reaching large enough amplitudes to produce a coherent and prominent peak in the frequency spectra. In all our simulations we found the presence of “critical layers”, regions where the drift frequency of the mode co-rotates with the local angular velocity of the flow. These regions are typically where non-axisymmetric inertial modes are expected to extract energy from the background shear flow and grow in amplitude (Rieutord2012). However, Baruteau2013 analyzed the interaction between inertial modes and differential rotation in a spherical shell, finding that for a cylindrical rotation profile of the form $\Omega(s)\propto 1+Cs^{2}$, modes are absorbed and damped at the co-rotation layer. Our simulations also produce cylindrical rotation, but with profiles far more complex than a simple quadratic in $s$, and the modes persist. Their persistence implies that either the simple profile result does not generalize, or another mechanism offsets corotation damping.
@@
The coherent inertial modes that are present appear to be linearly excited by a shear instability. Stochastic excitation seems unlikely because only a few modes are excited to large amplitudes in each simulation, whereas stochastic excitation typically excites many modes to similar amplitudes. In the Sun, inertial modes have recently been detected via helioseismology (see Gizon2024 and references therein), and some are excited stochastically (Philidet2023) while others are excited by a linear instability (Bekki2022b). Those authors attribute the excitation mechanism to baroclinic instability, but we believe this is unlikely in our case because our boundary conditions enforce nearly barotropic flow, as evidenced by the cylindrical flow patterns that develop. In any case, there is no obvious reason that similar mode excitation would not occur in real planets and stars.
@@
We have made several assumptions and approximations that should be relaxed in future work. For example, we adopted the Boussinesq approximation (Spiegel_Veronis_1960), which restricts the flow to be roughly incompressible, i.e., the density is approximately constant. Planets and stars have large density variations across their interiors and are highly compressible. As has been shown by several studies (LockitchFriedman1999; Wu2005; Mukhopadhyay2025), density stratification can change the power spectra at different radii, and affect the eigenfunctions of the modes, particularly those of non-toroidal modes with significant radial motions.
@@
In order to investigate inertial modes across different differential rotation profiles, we have focused our study on varying the convective Rossby number, $\mathrm{Ro_{c}}$, while keeping the Prandtl number fixed at unity. Yet, as shown in Section 3.5, lowering the Prandtl number to $\mathrm{Pr}=0.1$ leads to a far richer and more vigorous spectrum of inertial modes, with amplitudes exceeding those seen in the $\mathrm{Pr}=1$ cases. This suggests that the Prandtl number may play a more fundamental role in mode excitation than previously appreciated. Future simulations that push toward the astrophysically relevant regime $\mathrm{Pr}\ll 1$ would therefore be of great value for understanding inertial modes in planets and stars.
J.R.F. is supported by the Sherman Fairchild Postdoctoral (Burke) Fellowship and the Presidential Fellowship at Caltech, as well as NASA Solar System Workings grant 80NSSC24K0927. JF is grateful for support through the Caltech-JPL President’s and Director’s Research & Development Fund Program.

### v2

@@ III.2 Frequency Spectra
Figure 4 shows the power in the frequency spectrum of the flow at $r=0.85r_{o}$, for all the simulations with the exception of the non-rotating case, which is omitted as it shows no notable spectral features. Distinct coherent peaks in the spectrum reveal the excitation of modes, and the broader structure reflects the background of convective turbulence. The total frequency spectra refer to the ones summed over all $\ell$ and $m$, while ones for specific $m$ are summed over all $\ell$ for that specific order. Note that each spectrum contains an $m=0$ contribution that is much larger at very low frequencies. This corresponds to the “zero-frequency” mean flows associated with differential rotation.
@@
As the convective Rossby number decreases to 0.35 (panel d), a prominent $m=2$ mode appears, accompanied by a weaker $m=4$ mode whose frequency is twice that of $m=2$, indicating a non-linear self-interaction that resemble the ones described in Barik2018 for the spherical Couette system. At $\mathrm{Ro_{c}}\approx 0.26$ (panel e), the spectrum is dominated by an $m=3$ mode, followed by an $m=2$ signal at a nearby frequency, and low amplitude $m=1$ and $m=6$ modes. The frequency of the $m=6$ mode is exactly twice that of the $m=3$ mode, indicating that it arises from a self-interaction of the lower mode. At the lowest Rossby number in the suite, $\mathrm{Ro_{c}}\approx 0.17$ (panel f), we find three small-amplitude modes corresponding to $m=1$, $m=2$, and $m=6$, with the $m=2$ mode having a frequency roughly twice that of the $m=1$ mode. We emphasize that the amplitude of the excited modes depends on both the damping and the excitation mechanisms, which in our simulations are likely dominated by viscous dissipation and nonlinear interactions with convective motions (i.e. an effective turbulent viscosity), as well as by the degree of differential rotation. Although rotation weakens convection, particularly at low Rossby numbers, all of our rotating simulations exhibit strong differential rotation. However, we do not find any clear correlation between the mode amplitude and the strength of the differential rotation.
For comparison, each panel also shows the frequency of the spectrum normalized to the free-fall frequency $\omega_{\rm ff}\sim D/u_{\rm ff}$. We can see that the power spectrum is fairly flat for $\omega/\omega_{\rm ff}\lesssim 1$, and it generally falls steeply for $\omega/\omega_{\rm ff}\gtrsim 1$, as expected. However, it does not clearly exhibit the $\mathcal{P}(\omega)\propto\omega^{-2}$ spectrum expected for Kolmogorov turbulence with $\omega>\omega_{\rm ff}$ (goldreich:90; goldreich:94). We might expect different slopes in three regimes, with the lowest frequencies being $\omega<\omega_{\rm ff}$, the intermediate frequency regime in $\omega_{\rm ff}<\omega<2\Omega_{0}$, and the high-frequency regime being $\omega>2\Omega_{0}$. Our $\mathrm{Ro_{c}}\approx 0.17$ simulation appears to have a shallower slope to the power spectrum in the intermediate frequency regime than the high-frequency regime. For the other simulations, the separation in scales between $\omega_{\rm ff}$ and $2\Omega$ is too small to see any clear differences.
@@ III.3 Observed Wave Structures
Figure 6 shows the morphology of several modes excited in the simulations. We find that $u_{\phi}$ provides the strongest contribution to the power spectrum in these modes. Thus, we filter the zonal component of the velocity in space and time, isolating the azimuthal wavenumber $m$ and the mode frequency. The presence of differential rotation and contamination from convective motions makes it difficult to identify them in terms of pure eigenmodes. Nevertheless, these projections confirm that all the inertial modes in our simulations, except for the $m=0$ mode at $\mathrm{Ro_{c}}\approx 0.53$, exhibit symmetry about the equator, as also suggested by the dominance of the symmetric component in Figure 5.
III.4 Internal Shear Layers and Mode Attractors
Another useful measure is the meridional distribution of the zonal kinetic energy, which reveals the presence of internal shear layers (StewartsonRickard1969; Kerswell1995; Rieutord2001). These shear layers are aligned with the characteristic rays of inertial waves, which propagate at critical co-latitudes defined by $\cos\theta_{c}=\omega/2\Omega_{0}$, forming rays tilted by an angle $\pi/2-\theta_{c}$ with respect to the rotation axis. In a spherical shell, the corresponding boundary-value problem becomes ill-posed, producing singularities along these paths (StewartsonRickard1969; Kerswell1995). Viscosity removes these singularities and produces narrow regions of strong shear along the attractor paths.
@@ III.5 Low Prandtl number
We find several differences relative to the model using $\mathrm{Pr}=1$. For example, because the effective buoyancy scales as $\mathrm{Ra}/\mathrm{Pr}$, the flow velocities are much larger in the $\mathrm{Pr}=0.1$ simulation (compare the last two rows of Table 1). Another difference is the strength of both the convective flow and the differential rotation. While the $\mathrm{Pr}=1$ case shows weak flows at high latitudes (see panels a of Figures 1 and 2), the $\mathrm{Pr}=0.1$ model exhibits more vigorous flows at all latitudes (see the radial and azimuthal components of the velocity in Figure 8).
The frequency spectrum of the modes also shows remarkable differences with respect to the $\mathrm{Pr}=1$ case. As shown by Figure 9, mode excitation is more efficient at low Pr, with several large peaks appearing over the range m=1m=1–8. These inertial modes are mostly equatorially symmetric and drift retrograde in the rotating frame. In contrast, the $\mathrm{Pr}=1$ case exhibits much weaker excitation, with only three modes of very small amplitude (see panel f of Figure 4), even though the convective Rossby number $\mathrm{Ro_{c}}\approx 0.17$ remains the same in both simulations. As discussed previously, the dominant $m=1$ mode in this case is equatorially symmetric but prograde in the rotating frame. Finally, another difference is the spatial location of the modes. While the low amplitude $m=1$ mode observed for $\mathrm{Pr}=1$ is confined to the equatorial region (see Figure 6), the modes for $\mathrm{Pr}=0.1$ (including the $m=1$ mode) have most of their power from mid to high latitudes (see Figure 10). Though the onset of convection at low $\mathrm{Pr}$ occurs in the form of oscillatory quasi-geostrophic inertial-waves (QGIWs, Zhang2004), our simulation at $\mathrm{Pr}=0.1$ is sufficiently supercritical that we do not expect these QGIWs to play a significant role in the nonlinear dynamics. Nevertheless, a more systematic investigation of their influence will be addressed in future work.
IV Summary and Discussion
@@
In the spatial domain, the kinetic energy power spectra of the flow resemble those reported in previous numerical simulations of rotating convection (Featherstone2016; Hindman2020), with energy dominated by axisymmetric motions that produce strong peaks at odd spherical harmonic degrees $\ell$ (see Figure 3). Interestingly, the zonal component of the kinetic energy power spectra exhibits an $\ell^{-5}$ scaling, as observed in the spectra of the zonal flows of the gas and ice giants (e.g., Sanchez2000; Sukoriansky2002), argued for zonostrophic turbulence (see the recent preprint by Boning2024; Cabanes2024).
In the frequency domain, the low frequency range of the power spectra is dominated by the $m=0$ contribution of the differential rotation. At higher frequencies, the spectra show an overall flat distribution with superimposed peaks, followed by a rapid decay for $\omega/\omega_{\rm ff}>1$, i.e., for frequencies larger than the convective frequency (see Figures 4 and 9). These coherent peaks, which correspond to inertial modes excited within the flow, are only observed in cases with $\mathrm{Ro_{c}}\lesssim 0.53$, while they are absent at larger Rossby numbers. This is expected since the amount of differential rotation increases for more rapidly rotating flows. However, the amplitude of the peaks does not appear to correlate with the Rossby number and depends on the growth rate and saturation of each mode, influenced by both the damping and excitation mechanisms. We emphasize that the absence of inertial modes at larger Rossby numbers is specific to modes excited self-consistently by rotating convection in our setup, and may not hold for other excitation mechanisms or different choice of parameters. The Prandtl number nevertheless appears to play a more important role in mode excitation, since the $\mathrm{Pr}=0.1$ simulation exhibits a richer and more vigorous spectrum of inertial modes than any of the simulations with $\mathrm{Pr}=1$.
For the most part, the inertial modes in our simulations are nonaxisymmetric, retrograde in the rotating frame, symmetric about the equator, and span azimuthal orders $1\leq m\leq 8$, with frequencies $\omega/2\Omega_{0}\lesssim 0.2$ (see Figure 5 and 9). Their structures are complex, likely due to the effects of turbulent convection and differential rotation, and show that nearly all modes are confined to mid and high latitudes. This suggests that their excitation may be linked to shear instabilities within the differentially rotating convection zone. Further, their zonal energy distributions reveal internal shear layers that resemble mode attractors, particularly at high latitudes near the locations where the zonal flow changes sign (see Figure 7).
@@
Zhang2001 found a special class of equatorially symmetric inertial modes which strongly satisfy the Taylor-Proudman constraint (i.e., if the pressure gradient and the Coriolis force balance each other, see Proudman1916; Taylor1917). In such scenario, the modes are symmetric with respect to the equator, their frequencies satisfy $i\omega~{\mbox{{$u$}}}=-\nabla P/\rho-2\Omega_{0}{\mbox{{$\hat{z}$}}}\times{\mbox{{$u$}}}\approx 0$, and in turn, they are expected to be nearly geostrophic ($|\omega|/\Omega_{0}\ll 1$). Those modes are referred to as “slow” inertial modes (Barik2018). In contrast, other equatorially symmetric as well as antisymmetric modes that do not satisfy the Taylor–Proudman constraint can have much higher frequencies (see also Zhang2001; Wicht2014). The modes observed in our simulations are predominantly equatorially symmetric but have frequencies $\omega/2\Omega_{0}\sim 0.1$, suggesting that they are not close to the geostrophic limit.
Another difference is about the excitation mechanism. In general, the inertial modes observed in simulations and laboratory experiments of spherical Couette flows have been attributed to shear instabilities. It is likely that many of the modes in our simulations are excited simultaneously by both convection and differential rotation, with only a few reaching large enough amplitudes to produce a coherent and prominent peak in the frequency spectra. In all our simulations we found the presence of “critical layers”, regions where the drift frequency of the mode co-rotates with the local angular velocity of the flow. These regions are typically where non-axisymmetric inertial modes are expected to extract energy from the background shear flow and grow in amplitude (Rieutord2012; Astoul2021). However, Baruteau2013 analyzed the interaction between inertial modes and differential rotation in a spherical shell, finding that for a cylindrical rotation profile of the form $\Omega(s)\propto 1+Cs^{2}$, modes are absorbed and damped at the co-rotation layer. Our simulations also produce cylindrical rotation, but with profiles far more complex than a simple quadratic in $s$, and the modes persist. Their persistence implies that either the simple profile result does not generalize, or another mechanism offsets corotation damping.
@@
Although the amplitudes of the modes excited stochastically by convection depend on the frequency spectrum of the turbulent flow (goldreich:90), it typically leads to the excitation of a broad set of modes whose amplitudes reflect the underlying convective power spectrum. In contrast, in our simulations we observe only a small number of discrete inertial modes reaching large, long-lived, and phase-coherent amplitudes. This behavior is more naturally explained by linear excitation through a shear instability as seen in previous studies (Barik2018; Souza-Gomes2025). In the Sun, inertial modes have recently been detected via helioseismology (see Gizon2024 and references therein), and some are excited stochastically (Philidet2023) while others are excited by a linear instability (Bekki2022b). Those authors attribute the excitation mechanism to baroclinic instability, but we believe this is unlikely in our case because our boundary conditions enforce nearly barotropic flow, as evidenced by the cylindrical flow patterns that develop. In any case, there is no obvious reason that similar mode excitation would not occur in real planets and stars.
@@
We have made several assumptions and approximations that should be relaxed in future work. For example, we adopted the Boussinesq approximation (Spiegel_Veronis_1960), which restricts the flow to be roughly incompressible, i.e., the density is approximately constant. Planets and stars have large density variations across their interiors and are highly compressible. As has been shown by several studies (LockitchFriedman1999; Wu2005; Mukhopadhyay2025), density stratification can change the power spectra at different radii, and affect the eigenfunctions of the modes, particularly those of non-toroidal modes with significant radial motions. Density variations also change the morphology of the convective flow, creating a strong asymmetry with fast, narrow downflows and slow, wider upflows.
@@
In order to investigate inertial modes across different differential rotation profiles, we have focused our study on varying the convective Rossby number, $\mathrm{Ro_{c}}$, while keeping the Prandtl number fixed at unity. Yet, as shown in Section III.5, lowering the Prandtl number to $\mathrm{Pr}=0.1$ leads to a far richer and more vigorous spectrum of inertial modes, with amplitudes exceeding those seen in the $\mathrm{Pr}=1$ cases. This suggests that the Prandtl number may play a more fundamental role in mode excitation than previously appreciated. Future simulations that push toward the astrophysically relevant regime $\mathrm{Pr}\ll 1$ would therefore be of great value for understanding inertial modes in planets and stars.
We thank Adrian Barker for his careful review that helped us improve our manuscript. J.R.F. is supported by the Sherman Fairchild Postdoctoral (Burke) Fellowship and the Presidential Fellowship at Caltech, as well as NASA Solar System Workings grant 80NSSC24K0927. JF is grateful for support through the Caltech-JPL President’s and Director’s Research & Development Fund Program.